OpenStudy (gigi1103):

What is the equation, in point-slope form, for a line that goes through (8, −4)and has a slope of −5/6?

OpenStudy (gigi1103):

y−4=−5/6(x−8) y+4=−5/6(x+8) y−4=−5/6(x+8) y+4=−5/6(x−8)

OpenStudy (518nad):

equal of a line in point and slope form is if its point (x1,y1) and slope m then y-y1=m(x-x1) here u have point (8,-4), and slope -5/6

OpenStudy (518nad):

so which one is right?

OpenStudy (gigi1103):

a?

OpenStudy (3mar):

@518nad Don't leave her and go! You are her life saver!

OpenStudy (3mar):

Anyway, it would not be A. Think again please

OpenStudy (518nad):

i didnt leave her, she left me

OpenStudy (triciaal):